# Rethinking Model Prototyping through the MedMNIST+ Dataset Collection

### v2

@@ 3.4 Model Ranking
3.5 Quantitative Evaluation
In order to substantiate the qualitative findings presented earlier, we conduct a quantitative evaluation using non-parametric statistical tests. The aim is to discern the potential impact of input resolution (specifically, $28\times 28$, $64\times 64$, $128\times 128$, and $224\times 224$) and training scheme (comprising end-to-end training, linear probing, and $k$-NN integration) on the model performance, measured in terms of accuracy. Additionally, we aim to assess whether there are notable variations in performance across datasets depending on the model architecture. To this end, Friedman tests, with a significance level set at $p=0.05$, are initially employed to ascertain if statistically significant differences exist among the experimental conditions. Post-hoc two-tailed Wilcoxon signed-rank tests are then conducted to identify specific group differences, with a Bonferroni adjustment applied to mitigate the risk of type I errors resulting from multiple comparisons. Despite the limited number of observations ($n=10$ for model-related evaluations and $n=12$ for dataset-related analyses), we decided to assume an approximately normally distributed sampling distribution of the sample mean, following the Central Limit Theorem, in order to allow the computation of $Z$-values for the Wilcoxon tests. Finally, we asses the effect size of each test based on Cohen’s interpretation guidelines (Cohen, 1992) for Kendall’s Coefficient of Concordance $W$, and Pearson’s correlation coefficient $r$ for the Friedman and Wilcoxon tests, respectively.
@@ 3.5 Quantitative Evaluation
Next, we explore potential disparities in model performance based on the employed training schemes (end-to-end, linear probing, and $k$-NN). For this, we average the accuracy results per model and training scheme across all datasets and input resolutions. The Friedman test (with $dof=\text{number of training schemes}-1=2$) reveals overall significant differences suggesting a strong effect of the used training scheme on the model performance, $\left(\chi^{2}(\text{dof}=2)=14.6,\;p=0.0007,\;W=0.73\right)$. Furthermore, post-hoc Wilcoxon signed-rank tests with Bonferroni correction, adjusted significance level of $p<0.0167$, confirm these findings between end-to-end training and linear probing $\left(Z=-2.701,\;p=0.0039,\;r=0.854\right)$ as well as end-to-end training and $k$-NN $\left(Z=-2.803,\;p=0.002,\;r=0.886\right)$. These results, in conjunction with median IQR perceived accuracy values for end-to-end training of $82.77$ ($81.48$ to $83.61$), linear probing of $76.65$ ($74.45$ to $78.94$) and $k$-NN of $72.11$ ($72.37$ to $72.60$), respectively, affirm the superiority of end-to-end training in achieving the highest overall performance from Section 3.2. Interestingly, no significant differences are observed between linear probing and $k$-NN integration $\left(Z=-1.682,\;p=0.1055\right)$, despite a large effect size $\left(r=0.532\right)$. This suggests comparable efficacy of these training schemes despite variations in accuracy, thereby highlighting the potential of training-free strategies such as $k$-NN.
@@
Finally, we assess performance variations across datasets based on the model architecture. Through averaging accuracy results per dataset and model across all training schemes and input resolutions, we discern distinct performance trends. The IQR perceived accuracy for all models is visualized in Table 4 and the existence of statistical differences in dataset performance depending on the model architecture is illustrated in Table 5. Notably, DenseNet-121 and DINO ViT-B/16 emerged as the top-performing models, aligning with our previous findings of Section 3.4. Except of these, ViT-based models generally exhibit superior performance compared to convolutional models, which may seem contradictory to earlier findings. However, considering that we average the performance results across all training schemes, ViT models benefit from their higher performance for linear probing and $k$-NN integration compared to convolutional models, which only demonstrate higher performance in end-to-end training. Thus, these findings are indeed in line with our earlier observations. Furthermore, the results for SAM confirm its inferior performance compared to all other models as depicted in Section 3.2, thus contributing to a comprehensive understanding of model capabilities across various training contexts and tasks. Despite its similar performance for the end-to-end training, its task-foreign pretraining for segmentation rather than classification seems to limit its capabilities for training-less (linear probing) or training-free ($k$-NN) approaches.
@@
Further details regarding dataset-specific performance and model differences are elaborated in the appendix throughout Tables 6 to 17.
@@
4 Discussion and Conclusion
This work presents a comprehensive benchmarking analysis of convolutional and Transformer-based networks, for medical image classification across diverse datasets, training schemes, and input resolutions. Through systematic evaluation, we challenge prevailing assumptions regarding model design, training schemes, and input resolution requirements. Our experiments are designed to highlight both general dataset-average findings (see Table 3) and dataset-specific results (see Tables 6 - 17 in the appendix), which are basically coherent. By reassessing these methodologies, our aim is to foster genuine progress in the field and provide insights to inform the development of more efficient and effective models, rather than supporting the current trend of continuous scaling.
@@ 4 Discussion and Conclusion
However, it is crucial to acknowledge the limitations of our study. Our analysis is limited to the datasets provided by MedMNIST v2 and MedMNIST+ (Yang et al., 2023), and our results are dependent on their curation and distinct dataset splits. Furthermore, our evaluation across all datasets simultaneously may overlook dataset-specific nuances. Future studies should explore dataset-specific results, additional datasets with varying characteristics (i.e. sample size, noisy labels, corruptions, etc.), and different dataset splits. Additionally, the dataset collection lacks images from common modalities such as MRI, SPECT, and PET, and encompasses only a limited number of anatomical regions and disease patterns. Therefore, further investigation is warranted to include these modalities and assess the applicability of our findings in these unexplored settings. In conclusion, our work advocates for the following key takeaways and recommendations for model development:
@@
  • Prioritize the development of computationally efficient alternatives to end-to-end training for quicker model development iterations and reduced hardware strain during deployment.
@@
  • Consider utilizing lower resolution images during prototyping to conserve computational resources and time.
@@
  • Evaluate methods on multiple distinct benchmarks to cover real-world situations, rather than focusing solely on achieving state-of-the-art performance on a single benchmark.
@@
  • Focus on the development of efficient and robust methods, rather than scaling existing methods to attain state-of-the-art performance.
@@
Table 6: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the BloodMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 7: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the BreastMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 8: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the ChestMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 9: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the DermaMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 10: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the OctMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 11: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the OrganAMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 12: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the OrganCMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 13: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the OrganSMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 14: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the PathMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.

### v3

@@ Introduction
Quantitative Evaluation
In order to substantiate the qualitative findings presented earlier, we conduct a quantitative evaluation using non-parametric statistical tests. The aim is to discern the potential impact of input resolution (specifically, $28\times 28$, $64\times 64$, $128\times 128$, and $224\times 224$) and training scheme (comprising end-to-end training, linear probing, and $k$-NN integration) on the model performance, measured in terms of accuracy. Additionally, we aim to assess whether there are notable variations in performance across datasets depending on the model architecture. To this end, Friedman tests, with a significance level set at $p=0.05$, are initially employed to ascertain if statistically significant differences exist among the experimental conditions. Post-hoc two-tailed Wilcoxon signed-rank tests are then conducted to identify specific group differences, with a Bonferroni adjustment applied to mitigate the risk of type I errors resulting from multiple comparisons. Despite the limited number of observations ($n=10$ for model-related evaluations and $n=12$ for dataset-related analyses), we decided to assume an approximately normally distributed sampling distribution of the sample mean, following the Central Limit Theorem, in order to allow the computation of $Z$-values for the Wilcoxon tests. Finally, we asses the effect size of each test based on Cohen’s interpretation guidelines [48] for Kendall’s Coefficient of Concordance $W$, and Pearson’s correlation coefficient $r$ for the Friedman and Wilcoxon tests, respectively.
@@
Next, we explore potential disparities in model performance based on the employed training schemes (end-to-end, linear probing, and $k$-NN). For this, we average the accuracy results per model and training scheme across all datasets and input resolutions. The Friedman test (with $dof=\text{number of training schemes}-1=2$) reveals overall significant differences suggesting a strong effect of the used training scheme on the model performance, $\left(\chi^{2}(\text{dof}=2)=14.6,\;p=0.0007,\;W=0.73\right)$. Furthermore, post-hoc Wilcoxon signed-rank tests with Bonferroni correction, adjusted significance level of $p<0.0167$, confirm these findings between end-to-end training and linear probing $\left(Z=-2.701,\;p=0.0039,\;r=0.854\right)$ as well as end-to-end training and $k$-NN $\left(Z=-2.803,\;p=0.002,\;r=0.886\right)$. These results, in conjunction with median IQR perceived accuracy values for end-to-end training of $82.77$ ($81.48$ to $83.61$), linear probing of $76.65$ ($74.45$ to $78.94$) and $k$-NN of $72.11$ ($72.37$ to $72.60$), respectively, affirm the superiority of end-to-end training in achieving the highest overall performance from Section Benchmark Evaluation. Interestingly, no significant differences are observed between linear probing and $k$-NN integration $\left(Z=-1.682,\;p=0.1055\right)$, despite a large effect size $\left(r=0.532\right)$. This suggests comparable efficacy of these training schemes despite variations in accuracy, thereby highlighting the potential of training-free strategies such as $k$-NN.
@@
Finally, we assess performance variations across datasets based on the model architecture. Through averaging accuracy results per dataset and model across all training schemes and input resolutions, we discern distinct performance trends. The IQR perceived accuracy for all models is visualized in Table 3 and the existence of statistical differences in dataset performance depending on the model architecture is illustrated in Table 4. Notably, DenseNet-121 and DINO ViT-B/16 emerged as the top-performing models, aligning with our previous findings of Section Model Ranking. Except of these, ViT-based models generally exhibit superior performance compared to convolutional models, which may seem contradictory to earlier findings. However, considering that we average the performance results across all training schemes, ViT models benefit from their higher performance for linear probing and $k$-NN integration compared to convolutional models, which only demonstrate higher performance in end-to-end training. Thus, these findings are indeed in line with our earlier observations. Furthermore, the results for SAM confirm its inferior performance compared to all other models as depicted in Section Benchmark Evaluation, thus contributing to a comprehensive understanding of model capabilities across various training contexts and tasks. Despite its similar performance for the end-to-end training, its task-foreign pretraining for segmentation rather than classification seems to limit its capabilities for training-less (linear probing) or training-free ($k$-NN) approaches.
@@
Further details regarding dataset-specific performance and model differences are elaborated in the appendix throughout Supplementary Tables S1 to S12.
@@
Discussion and Conclusion
This work presents a comprehensive benchmarking analysis of convolutional and Transformer-based networks, for medical image classification across diverse datasets, training schemes, and input resolutions. Through systematic evaluation, we challenge prevailing assumptions regarding model design, training schemes, and input resolution requirements. Our experiments are designed to highlight both general dataset-average findings (see Table 2) and dataset-specific results (see Supplementary Tables S1 - S12 in the appendix), which are basically coherent. By reassessing these methodologies, our aim is to foster genuine progress in the field and provide insights to inform the development of more efficient and effective models, rather than supporting the current trend of continuous scaling.
@@
In addition to performance considerations, robustness to distribution shifts, model interpretability, and explainability are critical factors in clinical applications, where heterogeneous data is prevalent, and trust and transparency are paramount. While our work does not explicitly evaluate these aspects, prior research suggests that ViTs and ViT-based foundation models demonstrate greater robustness to distribution shifts compared to CNNs [49, 50, 51]. Furthermore, $k$-NN integration enhances interpretability compared to end-to-end training and linear probing by enabling direct, sample-based reasoning within the feature space [52, 53, 54]. Lastly, both CNNs and ViTs can be complemented with existing explainability techniques [55, 56], improving model transparency and aiding clinical decision-making. However, future work should further investigate these aspects to ensure that models not only achieve high performance but also provide clinically meaningful and interpretable insights for practitioners.
@@
However, it is crucial to acknowledge the limitations of our study. First, our analysis is limited to the datasets and distinct dataset splits provided by MedMNIST+ [33] making it inherently susceptible to biases such as class imbalance, data inhomogeneities, and demographic underrepresentation. In addition, the dataset collection lacks images from common modalities such as MRI, SPECT, and PET, and encompasses only a limited number of anatomical regions and disease patterns. Therefore, further investigation is warranted to include these modalities and assess the applicability of our findings in these unexplored settings. Second, our evaluation across all datasets simultaneously may overlook dataset-specific nuances. Future studies should explore dataset-specific results, additional datasets with varying characteristics (i.e. sample size, noisy labels, corruptions, etc.), and different dataset splits. Furthermore, our benchmark is restricted to Convolutional Neural Networks and Vision Transformer architectures. While these models remain fundamental to medical image classification, emerging architectures, such as Graph Neural Networks (GNNs), warrant further evaluation. GNNs have shown promise, particularly in histopathology, where they effectively capture topological structures in Whole Slide Images [57]. Expanding future benchmarks to include such architectures could provide a more comprehensive evaluation of the deep learning landscape for medical image classification. Additionally, we do not analyze the interpretability and explainability of the assessed architectures, which are critical for clinical adoption. Future work should address these aspects to enhance model transparency and trustworthiness. Finally, this work does not focus on the deployment of deep learning models in clinical practice which is a substantially more complex endeavor and requires addressing a broad set of challenges, including model explainability, data privacy, regulatory approval, and human-AI interaction. Instead, our benchmark aims to accelerate model development by providing insights that facilitate more efficient and scalable approaches for eventual real-world integration. In conclusion, our work advocates for the following key takeaways and recommendations for model development:
@@
  • Prioritize computational efficiency: prioritize the development of computationally efficient alternatives to full end-to-end training for faster model development iterations and reduced hardware demands during deployment.
@@
  • Utilize lower-resolution images for prototyping: consider utilizing lower resolution images during prototyping to conserve computational resources and time.
@@
  • Benchmark models across diverse datasets: evaluate methods on multiple distinct benchmarks to cover real-world situations, rather than focusing solely on achieving state-of-the-art performance on a single benchmark.
@@
  • Focus on efficiency and robustness: focus on the development of efficient and robust methods, rather than scaling existing methods to attain state-of-the-art performance.
Method
Model Selection
Our selection of model architectures encompasses a diverse array of both convolutional and Transformer-based networks. Among the chosen convolutional models are well-established architectures such as VGG16 [58], AlexNet [59], ResNet-18 [60], DenseNet-121 [61], and EfficientNet-B4 [62], all of which were pretrained on the ImageNet1k dataset [63]. In the domain of Transformers, we include the Vision Transformer (ViT) [3], renowned for its exceptional performance across various image classification tasks, pretrained on ImageNet1k as well as the CLIP [4] and DINO [5] pretrained ViT variants. Acknowledging recent advancements in this area, our selection extends to adaptations of ViT such as EVA-02 [64] and the Segment Anything Model (SAM) [30]. EVA-02 represents a series of efficiently optimized plain ViTs with moderate model sizes, employing bidirectional visual representations learned from a robust CLIP encoder. Conversely, SAM, initially conceived as a foundational model for image segmentation, is intentionally designed and trained to be promptable, thus facilitating zero-shot transferability to new image distributions and tasks, including image classification. For all ViT architectures, we opted for the base backbone with a patch size of 16 (i.e. ViT-B/16). With the exception of the AlexNet model obtained from the torchvision library, all models were sourced from the "Pytorch Image Models (timm)" library at Huggingface [65]. Further details regarding the employed backbone architectures, including the number of parameters, activations, Giga Multiply-Add Operations (GMACs), feature dimension before the final classification layer, memory requirements for single precision (fp32), and inference times on both GPU and CPU, are outlined in Table 5.
Table 5: Details of the evaluated model architectures, including the number of parameters (in million, M), activations (in million, M), Giga Multiply-Add Operations (GMACs), feature dimension before the final classification layer, memory requirements (in megabytes, MB) for single precision (fp32), and inference times on both GPU and CPU (in milliseconds, ms). GMACs, activations, inference times, and memory requirements are reported for processing a single input image of resolution 224 × 224 pixels. Inference times were measured on an Intel(R) Core(TM) i9-14900K processor (CPU) and an NVIDIA RTX 6000 GPU (Ada Generation).
Model Number Number Number Output Memory CPU GPU
Parameters Activations GMACs Dimension Requirements Inference Inference
(M) (M) # # (MB) (ms) (ms)
VGG16 138.4 13.6 15.5 4096 1454 108.0 0.4
AlexNet 62.3 0.6 0.36 4096 766 8.6 0.3
ResNet-18 11.7 2.5 1.8 512 602 9.2 0.7
DenseNet-121 8.0 6.9 2.9 1024 690 318.0 3.9
EfficientNet-B4 19.3 17.1 1.5 1792 728 29.3 3.9
ViT-B/16 86.6 16.5 16.9 768 900 49.0 1.3
CLIP ViT-B/16 86.6 16.5 16.9 768 900 48.9 1.3
EVA-02 ViT-B/16 86.3 16.5 16.9 768 902 60.0 2.8
DINO ViT-B/16 85.8 16.5 16.9 768 900 49.1 1.2
SAM ViT-B/16 89.7 64.3 23.3 256 924 71.4 13.2
Training Pipeline
We adopt diverse training paradigms, encompassing both end-to-end training (i.e. training the whole model), and linear probing, where we solely train the classification head, while keeping the encoder frozen. Additionally, we explore the integration of the k-nearest neighbors ($k$-NN) classifier into the feature space of pre-trained models. Following K. Nakata et al.[66] and S. Doerrich et al.[53], the pre-trained image encoder initially extracts feature embeddings from the training set, which are then stored in an external database along with their corresponding labels. During inference, the image encoder generates a feature embedding for a given query image. Subsequently, the top-$k$ feature embeddings having highest similarity scores (i.e. lowest cosine distance) with the query embedding are retrieved from the training set along with their associated labels. The classification of the query image is afterward determined through a majority vote on these labels. This approach facilitates efficient classification without necessitating retraining of the classification head or the entire encoder, thereby enhancing computational efficiency, interpretability, and generalizability, with reduced dependence on hyperparameters. Given the substantial computational cost associated with training deep neural networks, particularly foundation models, the adoption of $k$-NN methods presents an efficient alternative to traditional training schemes.
The training regimen consisted of 100 epochs with early stopping based on the validation set. We employed the AdamW optimizer [67] with a learning rate of $0.0001$, along with a cosine annealing learning rate scheduler [68] with a single cycle. Each model was trained with a batch size of $64$, allowing for training on a single NVIDIA RTX™ A5000 GPU. For evaluations utilizing the k-nearest neighbors ($k$-NN) approach, we set $k$ to $11$, in line with Z. Zhu et al. [69], who demonstrated the suitability of $k>10$ for detecting noisy labels. To maintain compatibility with the pretrained models while preserving the inherent properties of individual resolutions, all image resolutions were padded to $224\times 224$ pixels using zero padding. This choice was motivated by M. Hashemi [70], who demonstrated that zero-padding has no discernible effect on classification accuracy while significantly reducing training time compared to image resizing. With zero-padding, neighboring zero input units (pixels) do not activate their corresponding convolutional unit in the subsequent layer, resulting in decreased requirements for updating synaptic weights on outgoing links and ensuring robust feature preservation during image reshaping.
Loss Criterion and Evaluation Metrics
In line with the methodology outlined by J. Yang et al. [32], we select the choice of loss criteria to suit the specific classification tasks associated with each dataset. For binary (BC) and multi-class classification (MC), as well as ordinal regression (OR) tasks, we utilize the Cross-Entropy (CE) loss function applied to the logits:
Furthermore, for the purpose of simplicity and standardization, consistent with J. Yang et al. [32], we employ the same evaluation metrics including accuracy (ACC) for a fixed operating point of $0.5$ and the area under the receiver operating characteristic curve (AUC) to assess the model’s ability to differentiate between classes. Our choice of AUC over metrics such as sensitivity, specificity, and false positive rates is driven by its suitability for comparing different architectural choices and training paradigms. Unlike sensitivity and specificity, which require setting a fixed operating point, a process that is application-dependent and not trivial, AUC provides a holistic measure of a model’s discriminatory power across all operating points. Additionally, we include accuracy due to its prevalence in the field and ease of interpretation.
To quantify the level of class imbalance within each dataset, we report Shannon’s Equitability for each data split. Shannon’s Equitability (EH) is a normalized measure of class distribution uniformity, derived from the Shannon Diversity Index [71]. It is computed as:
$EH=\frac{H}{H_{\text{max}}}=\frac{-\sum_{i=1}^{S}p_{i}\ln{p_{i}}}{\ln{S}},$ (3)
where $H$ is the Shannon Diversity Index, $S$ represents the total number of classes, $p_{i}$ denotes the proportion of samples belonging to class $i$ and $\ln$ stands for the natural logarithm. The maximum possible diversity, $H_{\text{max}}=\ln{S}$, occurs when all classes are equally represented. Shannon’s Equitability ranges from 0 (total imbalance) to 1 (perfect balance), allowing for a standardized comparison of class distribution across datasets and splits.
Data availability
The used datasets are licensed under Creative Commons Attribution 4.0 International (CC BY 4.0), except DermaMNIST under Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) and publicly available here: Yang, J. et al. [MedMNIST+] 18x Standardized Datasets for 2D and 3D Biomedical Image Classification with Multiple Size Options: 28 (MNIST-Like), 64, 128, and 224. https://zenodo.org/records/10519652 (2024).
Code availability
Acknowledgements
Funded through the Hightech Agenda Bayern (HTA) of the Free State of Bavaria, Germany.
Author contributions statement
S.D., F.D.S., J.B., and C.L. worked on the methods and did the study design. S.D., F.D.S., and J.B. implemented the algorithms. J.B. analyzed the individual data sets in detail. S.D. and F.D.S. conceived and conducted the experiments. S.D. analyzed the results and created all images. C.L. supervised the study. S.D., F.D.S., and C.L. wrote the manuscript. All authors reviewed, corrected, and approved the paper.
Additional information
Compliance with ethical standards: This research study was conducted retrospectively using human subject data made available in open access. Ethical approval was not required as confirmed by the license attached with the open-access data.
Competing interests: The authors declare no competing interests.
@@ Acknowledgements
Supplementary Table S1: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$, and area under the receiver operating characteristic curve (AUC) for the BloodMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S2: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$, and area under the receiver operating characteristic curve (AUC) for the BreastMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S3: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the ChestMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S4: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the DermaMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S5: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the OctMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S6: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the OrganAMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S7: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the OrganCMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S8: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the OrganSMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S9: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the PathMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.